# Scalable Generalized Bayesian Online Neural Network Training for Sequential Decision Making

### v1

@@ C.2 Singular value decomposition given sum of low-rank-matrices
Hence $\sum_{i}\mathbf{A}_{i}$ has eigenvectors $\mathbf{V}$ and eigenvalues given by the diagonal entries of $\mathbf{S}^{2}$. By the symmetric-matrix form of the Eckart–Young theorem [19, 12], its best rank-$d$ approximation in Frobenius norm is
@@
$\mathbf{S}_{:d}$ $:={\rm diag}\left(\sigma_{1},\dots,\sigma_{d}\right),$
@@
$\mathbf{S}_{:d,+q}$ $:={\rm diag}\left(\sqrt{\sigma_{1}^{2}+q},\dots,\sqrt{\sigma_{d}^{2}+q}\right),$
@@
Appendix D Further results on our method
D.1 Derivation
@@
We start the algorithm by initialising the beliefs about the last layer parameters ${\boldsymbol{\ell}}$ and the hidden layer parameters ${\bm{h}}$. We set
@@
for known $\mathbf{C}_{0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$, ${\bm{h}}_{0|0}\in\mathbb{R}^{D_{{\bm{h}}}}$, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$, and ${\boldsymbol{\ell}}_{0|0}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$. As mentioned above, the notation ${\bm{\theta}}_{t|k}$ (for $k,t\in\mathbb{N}$) stands for the best linear unbiased predictor (BLUP) of ${\bm{\theta}}_{t}$ given the information up to time $k$. Next, we assume that for $t\geq 1$, the latent last layer parameters ${\boldsymbol{\ell}}_{1:t}$ and the hidden layer parameters ${\bm{h}}_{1:t}$ follow the dynamics
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\$ (43)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables333 i.e., ${\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},i},{\bm{u}}_{{\bm{h}},j})=\boldsymbol{0}$ for all $i,j\in\{1,\ldots,T\}$. with dynamics covariance matrices ${\rm Var}({\bm{u}}_{{\boldsymbol{\ell}},t})=\mathbf{Q}_{{\boldsymbol{\ell}},t}\in{\mathcal{M}^{+}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ and ${\rm Var}({\bm{u}}_{{\bm{h}},t})=\mathbf{Q}_{{\bm{h}},t}\in{\mathcal{M}^{+}_{{D_{{\bm{h}}}}}(\mathbb{R})}$. For simplicity and to exploit efficient linear algebra techniques for belief updates, we assume $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$. A simple choice is to set $\mathbf{Q}_{{\bm{h}},t}=0\mathbf{I}$ to avoid forgetting past data; see e.g., [18].
@@
The proof of Proposition D.1 is in Appendix F.1.
@@
The proof of Proposition D.2 is in Appendix F.2.
@@ D.2 Computational complexity
Appendix E Algorithms
E.1 HiLoFi
See section D.1 for the derivation and algorithm 2 for the code.
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $(q_{{\boldsymbol{\ell}},t},q_{{\bm{h}},t})$ dynamics covariance for last layer and hidden layers
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $\left({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\mathbf{C}_{t-1}\right)$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$
3:  $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
4:  $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
5:  // innovation (one-step-ahead error) and Cholesky innovation variance
6:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
7:  $\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$
8:  // gain for hidden layers
9:  $\mathbf{V}_{{\bm{h}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{H}}_{t}$
10:  $\mathbf{K}_{{\bm{h}},t}^{\intercal}\leftarrow\mathbf{V}_{{\bm{h}},t}\,\mathbf{C}_{t-1}\,\mathbf{C}_{t-1}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{V}_{{\bm{h}},t}$
11:  // gain for last layer
12:  $\mathbf{V}_{{\boldsymbol{\ell}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{L}}_{t}$
13:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
14:  // BLUP update step
15:  ${\bm{h}}_{t|t}\leftarrow{\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\boldsymbol{\epsilon}_{t}$
16:  ${\boldsymbol{\ell}}_{t|t}\leftarrow{\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\boldsymbol{\epsilon}_{t}$
17:  // EVC low-rank and Cholesky update step
18:  $\mathbf{C}_{t}\leftarrow{\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)$
19:  $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\leftarrow{\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right)$
20:  return  $\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\right)$
Algorithm 2 Single step of HiLoFi for online learning.
E.2 LoLoFi
@@
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $(q_{{\boldsymbol{\ell}},t},q_{{\bm{h}},t})$ dynamics covariance for last layer and hidden layers
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $\left({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},\mathbf{C}_{{\boldsymbol{\ell}},t-1},\mathbf{C}_{{\bm{h}},t-1}\right)$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$
3:  $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
4:  $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
5:  // innovation (one-step-ahead error) and Cholesky innovation variance
6:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
7:  $\mathbf{S}_{t}^{1/2}\leftarrow{\cal Q}_{R}\left({\mathbf{C}}_{{\boldsymbol{\ell}},t-1}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{{\bm{h}},t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$
8:  // gain for hidden layers
9:  $\mathbf{V}_{{\bm{h}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{H}}_{t}$
10:  $\mathbf{K}_{{\bm{h}},t}^{\intercal}\leftarrow\mathbf{V}_{{\bm{h}},t}\,\mathbf{C}_{{\bm{h}},t-1}\,\mathbf{C}_{{\bm{h}},t-1}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{V}_{{\bm{h}},t}$
11:  // gain for last layer
12:  $\mathbf{V}_{{\boldsymbol{\ell}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{L}}_{t}$
13:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
14:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\mathbf{C}_{{\boldsymbol{\ell}},t-1}\,\mathbf{C}_{{\boldsymbol{\ell}},t-1}^{\intercal}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
15:  // mean update step
16:  ${\bm{h}}_{t|t}\leftarrow{\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\boldsymbol{\epsilon}_{t}$
17:  ${\boldsymbol{\ell}}_{t|t}\leftarrow{\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\boldsymbol{\epsilon}_{t}$
18:  // low-rank updates
19:  ${\mathbf{C}}_{{\boldsymbol{\ell}},t}\leftarrow{\cal P}_{{d_{{\boldsymbol{\ell}}}},+q_{{\boldsymbol{\ell}},t}}\left(\mathbf{C}_{{\boldsymbol{\ell}},t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right)$
20:  ${\mathbf{C}}_{{\bm{h}},t}\leftarrow{\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{{\bm{h}},t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)$
21:  return  $\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{{\boldsymbol{\ell}},t},\mathbf{C}_{{\bm{h}},t}\right)$
@@
E.3 Low-rank Kalman filter (LRKF)
@@
Proposition E.1 (Kalman gain and innovations).
@@
Proposition E.2 (Update step).
@@
Next, we consider the surrogate covariance
@@
$\hat{\boldsymbol{\Sigma}}_{t}$ $=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t|t}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$ (65)
@@
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $q_{t}$ // dynamics covariance
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $(\boldsymbol{\mu}_{t-1},\mathbf{C}_{t-1})$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow h({\bm{\theta}}_{t},{\bm{x}}_{t})$
3:  $\mathbf{H}_{t}\leftarrow\nabla_{\bm{\theta}}h(\boldsymbol{\mu}_{t-1},{\bm{x}}_{t})$
4:  // innovation and (Cholesky) innovation variance
5:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
6:  $\mathbf{S}_{t}^{1/2}\leftarrow{\cal Q}_{R}(\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal},\sqrt{q_{t}}\,\mathbf{H}_{t},\,\mathbf{R}_{t}^{1/2})$
7:  // gain matrix
8:  $\mathbf{V}_{t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\,\mathbf{H}_{t}$
9:  $\mathbf{K}_{t}^{\intercal}\leftarrow\mathbf{V}_{t}\,\mathbf{C}_{t-1}\,\mathbf{C}_{t-1}^{\intercal}+q_{t}\,\mathbf{V}_{t}$ // low-rank update
10:  // mean and low-rank factor updates
11:  $\boldsymbol{\mu}_{t}\leftarrow\boldsymbol{\mu}_{t-1}+\mathbf{K}_{t}\,\boldsymbol{\epsilon}_{t}$
12:  $\mathbf{C}_{t}\leftarrow{\cal P}_{d}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\right)$
13:  return  $(\boldsymbol{\mu}_{t},\mathbf{C}_{t})$

### v2

@@ C.2 Singular value decomposition given sum of low-rank-matrices
Hence $\sum_{i}\mathbf{A}_{i}$ has eigenvectors $\mathbf{V}$ and eigenvalues given by the diagonal entries of $\mathbf{S}^{2}$. By the symmetric‐matrix form of the Eckart–Young theorem [17, 10], its best rank-$d$ approximation in Frobenius norm is
@@
$\mathbf{S}_{:d}$ $={\rm diag}\left(\sigma_{1},\dots,\sigma_{d}\right),$
@@
$\mathbf{S}_{:d,+q}$ $={\rm diag}\left(\sqrt{\sigma_{1}^{2}+q},\dots,\sqrt{\sigma_{d}^{2}+q}\right),$
@@
Appendix D HiLoFi— further details
D.1 Derivation of HiLoFi
@@
[…]
@@
for known $\mathbf{C}_{0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$, ${\bm{h}}_{0|0}\in\mathbb{R}^{D_{{\bm{h}}}}$, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$, and ${\boldsymbol{\ell}}_{0|0}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$. Next, we assume that for $t\geq 1$, the latent last layer parameters ${\boldsymbol{\ell}}_{1:t}$ and the hidden layer parameters ${\bm{h}}_{1:t}$ follow the dynamics
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\\$ (48)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables111i.e., ${\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},i},{\bm{u}}_{{\bm{h}},j})=\boldsymbol{0}$ for all $i,j\in\{1,\ldots,T\}$. with dynamics covariance matrices ${\rm Var}({\bm{u}}_{{\boldsymbol{\ell}},t})=\mathbf{Q}_{{\boldsymbol{\ell}},t}\in{\mathcal{M}^{+}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ and ${\rm Var}({\bm{u}}_{{\bm{h}},t})=\mathbf{Q}_{{\bm{h}},t}\in{\mathcal{M}^{+}_{{D_{{\bm{h}}}}}(\mathbb{R})}$. For simplicity and to exploit efficient linear algebra techniques for belief updates, we assume $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$. A simple choice is to set $\mathbf{Q}_{{\bm{h}},t}=0\mathbf{I}$ to avoid forgetting past data; see e.g., [16].
@@
The proof of Proposition D.1 is in Appendix E.1.
@@
The proof of Proposition D.2 is in Appendix E.2.
@@ D.2 Computational complexity
Appendix E Proofs
E.1 Proof of Proposition D.1
Proof.
Following Proposition B.2, the predicted mean takes the form
Next, the predicted posterior covariance takes the form
where
$\boldsymbol{\Sigma}_{t-1|t-1}$ (57)
$\quad={\rm Var}({\bm{\theta}}_{t-1}-{\bm{\theta}}_{t-1|t-1})$
Next,
$\quad={\rm Cov}({\boldsymbol{\ell}}_{t-1},{\bm{h}}_{t-1})$
$\quad=\boldsymbol{0},$
which follows from the assumptions about the SSM. Then
and
Given that $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}$ is approximated at t1t-1 through $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}$ and $\boldsymbol{\Sigma}_{{\bm{h}},t-1|t-1}$ is approximated at t1t-1 through $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}$, we obtain
E.2 Proof of Proposition D.2
Proof.
$=\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1})+{\bm{e}}_{t}$
Next, the variance of the innovation is
$\mathbf{S}_{t}$ $={\rm Var}(\boldsymbol{\epsilon}_{t})$ (63)
$=\mathbf{S}_{t}^{\intercal/2}\,\mathbf{S}_{t}^{1/2},$
with
E.3 Proof of Proposition D.3
Proof.
First, by Proposition B.2, the Kalman gain matrix is given by
where
$\mathbf{K}_{{\boldsymbol{\ell}},t}$ $=\Big(\mathbf{L}_{t}^{-1}\,\mathbf{L}_{t}^{-\intercal}\tilde{\mathbf{L}}_{t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\Big)^{\intercal},$ (67)
where
$\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t}$ (69)
the predicted error covariance $\boldsymbol{\Sigma}_{t|t-1}$ is
and
Thus,
$\boldsymbol{\Sigma}_{t|t}$ (72)
Next, we consider the following surrogate covariance matrix which modifies the second block-diagonal entry
$\tilde{\boldsymbol{\Sigma}}_{t|t}=$ (73)
$\left[\begin{smallmatrix}\tiny(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \tiny\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}\end{smallmatrix}\right].$
It then follows that
with
Next, we find the update rule for the Cholesky and low-rank factors. First, the EVC for the last layer takes the form
so that
Then, the EVC covariance for the hidden layers (lower right block-diagonal) is
Finally, by Corollary C.3, the best rank-$d$ low-rank factor is
So that
E.4 Proof of Proposition 4.1
Proof.
We seek to bound
Following the arguments from Proposition F.3, we obtain
where
and
Appendix F Further algorithms
F.1 LoLoFi
@@
[…]
5:// predict step
[…]
9:// innovation(one-step-ahead error) and Cholesky innovation variance
[…]
12:// gain for hidden layers
[…]
15:// gain for last layer
[…]
19:// mean update step
[…]
22:// low-rank updates
[…]
25:Return ${\bm{b}}_{t}=\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{{\boldsymbol{\ell}},t},\mathbf{C}_{{\bm{h}},t}\right)$
@@
F.2 Low-rank Kalman filter (LRKF)
@@
Proposition F.1 (Kalman gain and innovations).
@@
Proposition F.2 (Update step).
@@
Next, we consider the surrogate EVC matrix
@@
$\hat{\boldsymbol{\Sigma}}_{t}$ $=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$ (101)
@@
[…]
5:// predict step
[…]
8:// innovation and (Cholesky) innovation variance
[…]
11:// gain matrix
[…]
14:// mean and low-rank factor updates
[…]
17:Return ${\bm{b}}_{t}=(\boldsymbol{\mu}_{t},\mathbf{C}_{t})$ // updated belief